How much water should be added to 10 liters of a 45% alcohol solution to produce a 30% solution?
Answer by stanbon(75887)

How much water should be added to 10 liters of a 45% alcohol solution to produce a 30% solution?
-------
Equation:
water + water = water
0.55*10 + 1*x = 0.70(10+x)
------------------
55*10 + 100x = 70*10 + 70x
30x = 15*10
x = 5 liters (amt. of water to add)
====================
Cheers,
Stan H.
====================